Please draw each point

PROBLEM 4 (25%) Given the following state of stress, 20 ksi 16 ksi 8 ksi Given the following state of stress, using the MohrC (ksi)y PRINCIPAL STRESS (ksi) Ox + y = ksi y X MAX SHEAR STRESS (ksi) Ox + Oy ksi STRESS @ 35 deg CCW (ksi) Ox + Oy = ksi
Answer #1

20 ksi stress element *8 ksi 8 ksi 16 ks, 10. 20 ksi Coordinate of ** Face of stress element, (5x, exy) (8,16) A B = ( Oy, TNormal to x 35 x 2 = 70° 3x2-x of pulona Emay = 21.26 616.45,18.63) Ri Al 8,16) 41.189 (20p) 48.820 max. principal stress 0 :Resulty from mohrle circle 5,= max, principal stress o+R Radius - COA) (3%-(-6)) + (06-0}} 0,- -6+21.26 R 2 1.26 ksi 15,= 15,stress element 2 stress in varience 15.26 —27.2. = , 1 1-2ч.ч)° 0 : 15.24 6. : 27. 24 6, + бу - 8+(-20) = - 2 15.24 +(-27, 24Maximum shear stress = Tmax=+R = + 21.26 KSI Tomax e + 21,26 ksi maximum shear stress normal stress associated with plane -6kGrigel 2 р 20.26 1с Д...24 -21,2 6 6, +6y = 8-20 = Cl2 644 ! – – – 6 — -12 Srix+ Smy б, toy = иелло Onx tony20p 4 20s ) (31) tant () 200= 이 • 4 3.82 | 20 , 4 2-2u.u CI) AG 20 2. 2092 180 + 207) 아6) - 228.32 Or, < 114.1°C 20 A(8,16) Ucoordinate of Rotated plane (R2gR 2) (0,,7,) (RJI O 70 R = 21,26 61.18 48.82 (1.80 66,0) (-60) 에 sin (61.18) o = cos( 81.1coordinate of Rotated plane (R2gR 2) (0,,7,) (RJI O 70 R = 21,26 61.18 48.82 (1.80 66,0) (-60) 에 sin (61.18) o = cos( 81.1(0 م ) 5- 5.26 ks (1.176 R = 21,28 « ( 6, 7) Tmax وا (61.125)، ما Rx + 6 - 16, 9. 24 ks : ج R sin( 600) 89) - 21.2 x sin (6stress element * stress invarionce Б. си Ч,ч - 4,25 Byla 350 1861 - 12.01 Ч,241 : 2-20 бу б) +бу 6. - 16,25 -2 -16,25 +4,247